# Predicting the outcome of collisional neutrino flavor conversion

### v2

@@ Abstract
Collisional flavor instabilities, driven by differing neutrino and antineutrino reaction rates, are expected to occur in dense astrophysical environments like supernovae and neutron star mergers, but have yet to be incorporated in large-scale simulations. We derive analytical expressions for the asymptotic state resulting from a homogeneous and isotropic instability, and apply these predictions to two representative conditions from a neutron star merger simulation. We emphasize the importance of using a collision term that allows for both damping of flavor coherence and relaxation back to the classical steady-state. When this classical configuration is collisional-unstable, the resulting asymptotic state reflects a compromise between classical relaxation and flavor conversion, defining a “quantum” equilibrium with nonzero coherence. This analysis highlights the possibility of a trade-off between classical and quantum effects, an important feature with regard to the inclusion of flavor oscillation physics into global simulations.
@@ I Introduction
In our current era of multi-messenger astronomy, significant efforts are underway to model dense astrophysical objects such as core-collapse supernovae (CCSNe) or binary neutron star mergers (NSMs), with large-scale simulations incorporating a wide range of physics, including (magneto)hydrodynamics, general relativity, or the nuclear equation of state (see [1, 2, 3, 4, 5, 6] and references therein). A key ingredient in all simulations is neutrino transport [7, 8, 9, 10] yet to date, neutrino oscillations are not systematically taken into account. Over the last decades, however, various flavor transformation mechanisms have been identified, which are expected to occur in these environments and may have important consequences on the dynamics (see, e.g., [11, 12, 13, 14, 15, 16, 17, 18] for reviews).
@@
A direct approach, solving the “quantum” version of Boltzmann’s equation describing neutrino transport with flavor mixing—the Quantum Kinetic Equation (QKE)—, is made unfeasible by the vast differences of length- and time-scales involved. Nevertheless, to assess the potential consequences of neutrino flavor oscillations, simplified models have been used in CCSNe [19, 20, 21, 22], post-merger accretion disks [23, 24, 25] and recently a full NSM simulation [26]. Other techniques, designed to mitigate the computational cost of a full multi-angle QKE treatment, include an attenuation of the QKE Hamiltonian [27, 28, 29, 30, 31, 32], or the use of angular moments of the neutrino density matrix [33, 34, 35, 36, 37, 38, 39]. A convenient approach that can be embedded in classical simulations consists in designing a subgrid model of flavor transformation [40], which requires one to determine the state reached by the system post-flavor conversion. This has been heavily studied for fast flavor instabilities (FFIs), establishing that the final quasi-steady state is characterized by some cancellation of the angular crossing between (anti)neutrino distributions, which can be predicted in various ways [41, 42, 43, 44, 45, 46, 47, 48, 49]. Such prescriptions have recently been used in global simulations of a post-merger disk [50] and CCSNe [51].
@@
A similar analytical prediction of the flavor conversion outcome remains, so far, unrealized for “collisional” flavor instabilities (CFIs), first evidenced in [52]. CFIs take place as neutrino self-interactions combine with different neutrino/antineutrino reaction rates to amplify flavor coherence, contrary to the standard expectation that collisions damp flavor coherence. Linear stability analysis [53, 54, 55, 56, 57] for a homogeneous and isotropic background has shown the existence of different collisional-unstable modes, with resonance-like and non-resonance-like regimes. The occurrence of CFIs has been investigated in actual environments, namely CCSNe [28, 58, 59, 60, 61, 62] and NSMs [53, 36, 63]. However, most studies consider a simplified collision term which only acts on flavor coherence. This still leads to a rich phenomenology, notably with multi-energy neutrino gases, from flavor equipartition to flavor swaps [54, 55, 64, 65, 66]. Including the flavor-diagonal parts of the collision term is nevertheless important, as they lead for instance to a flavor swap in the resonance-like case [67], instead of flavor equipartition.
@@
This paper is structured as follows: in Sec. II, we present our model of neutrino evolution, which includes both classical relaxation and coherence damping components of the collision term. We also present the associated linear stability results, with details in Appendix A. In Sec. III, we numerically solve the evolution equations for configurations taken from a classical NSM simulation (see Appendix B for a convergence study), and derive analytical predictions of the asymptotic state that agree with the non-linear outcome of collisional flavor conversion. These results show significant differences with earlier studies that use different implementations of the collision term, which is examined in Sec. IV. We summarize our findings and discuss the limitations and prospects of this study in Sec. V.
@@ II.1 Model and assumptions
where $\omega\equiv\Delta m^{2}/2E_{\nu}$ with $E_{\nu}$ the common neutrino energy, and we use the compact notation $c_{2\theta}=\cos(2\theta)$ and $s_{2\theta}=\sin(2\theta)$, with $\theta$ the mixing angle. In this work, we take $\Delta m^{2}=2.5\times 10^{-3}\,\mathrm{eV^{2}}$ and $\sin^{2}(\theta)=0.307$ (values corresponding to $\Delta m^{2}_{12}$ and $\theta_{12}$ in the 3-flavor case [71]). The matter term reads, also in the flavor basis,
@@
The last terms in Eq. (1) are an approximate description of collisions, using a “relaxation-time” approximation. The quantities $N^{(\mathrm{cl})}$, $\overline{N}^{(\mathrm{cl})}$ are flavor-diagonal and describe the classical steady-state of the system. The timescales associated to this relaxation are given by the opacities $\kappa_{e}$, $\bar{\kappa}_{e}$ and $\kappa_{x}=\bar{\kappa}_{x}$ (we assume, consistently with many large-scale simulations, identical evolutions for $\nu_{x}$ and $\bar{\nu}_{x}$). For $\nu_{e}$ and $\bar{\nu}_{e}$, this corresponds to emission/absorption processes, with absorption opacity $\kappa_{e}$ and associated emissivity $\kappa_{e}N_{ee}^{(\mathrm{cl})}$ (likewise for $\bar{\nu}_{e}$), but the treatment of heavy-lepton flavor (anti)neutrinos is more subtle. Indeed, it was shown in [61] that a system in beta- and pair-equilibrium is stable to CFIs. In an actual system that is unstable to CFI, the heavy-lepton flavor neutrinos $\nu_{x}$ deviate from this thermal equilibrium as they are not perfectly trapped. The classical steady-state $N_{xx}^{(\mathrm{cl})}$ is thus the result of a competition between emission/absorption and advection. An effective homogeneous description of the system can be done in a relaxation-time approximation, assuming the same form of a collision-like term as for electron flavor (anti)neutrinos but with an effectively reduced emissivity (see Sec. III.1 for an actual example), which leads to the form (1) of the evolution equation. In the following, we conflate “collisions” with “classical steady-state relaxation,” bearing in mind that this remains a limitation of our homogeneous description. We note that this description would also apply to a perfectly homogeneous and isotropic system with a non-standard classical equilibrium. Finally, since we consider isotropic neutrino distributions, the scattering processes can be discarded.
We will assume that initially, the system is in the classical steady-state $N(t=0)=N^{(\mathrm{cl})}\equiv\mathrm{diag}\left(N_{ee}^{(\mathrm{cl})},N_{xx}^{(\mathrm{cl})}\right)$, and likewise for antineutrinos. This assumption allows us to focus on the competing effects of flavor conversion and classical relaxation. If the initial configuration was not a classical steady-state, the diagonal components would evolve on the collisional timescale while flavor coherence grows, but accounting for these effects would notably require to describe the advection of $\nu_{x}$. We thus focus on an idealized situation which allow for an analytical treatment, and provides guidance for future studies on larger spatial scales.
@@ II.2 Linear regime
We first establish the main features of the CFI in the linear regime, where $N_{aa}\simeq N_{aa}^{(\mathrm{cl})}$ and $\lvert N_{ex}\rvert\ll 1$. Following [54], we write the $ex$ components of the QKEs as
$\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}N_{ex}\\ \overline{N}_{ex}\end{bmatrix}\simeq\frac{\omega s_{2\theta}}{2\sqrt{2}G_{F}}\begin{bmatrix}-\mathfrak{g}^{(\mathrm{cl})}\\ \bar{\mathfrak{g}}^{(\mathrm{cl})}\end{bmatrix}+\Lambda\begin{bmatrix}N_{ex}\\ \overline{N}_{ex}\end{bmatrix}\,,$ (4)
@@
$\Lambda=\begin{bmatrix}-\omega c_{2\theta}+\lambda-\bar{\mathfrak{g}}^{(\mathrm{cl})}-\mathrm{i}\Gamma&\mathfrak{g}^{(\mathrm{cl})}\\ -\bar{\mathfrak{g}}^{(\mathrm{cl})}&\omega c_{2\theta}+\lambda+\mathfrak{g}^{(\mathrm{cl})}-\mathrm{i}\overline{\Gamma}\end{bmatrix}\,,$ (5)
@@
$\mathrm{Im}\left(\Omega_{-}\right)$ $=\frac{\overline{\Gamma}\bar{\mathfrak{g}}^{(\mathrm{cl})}-\Gamma\mathfrak{g}^{(\mathrm{cl})}}{\mathfrak{g}^{(\mathrm{cl})}-\bar{\mathfrak{g}}^{(\mathrm{cl})}}\,,$ $Q_{-}$ $\propto\begin{bmatrix}\mathfrak{g}^{(\mathrm{cl})}\\ \bar{\mathfrak{g}}^{(\mathrm{cl})}\end{bmatrix}\,,$ (7a)
$\mathrm{Im}\left(\Omega_{+}\right)$ $=\frac{\Gamma\bar{\mathfrak{g}}^{(\mathrm{cl})}-\overline{\Gamma}\mathfrak{g}^{(\mathrm{cl})}}{\mathfrak{g}^{(\mathrm{cl})}-\bar{\mathfrak{g}}^{(\mathrm{cl})}}\,,$ $Q_{+}$ $\propto\begin{bmatrix}1\\ 1\end{bmatrix}\,.$ (7b)
We assumed that we are outside the resonance-like regime, which is the case in the examples studied below.
@@ III.1 NSM examples
We consider two configurations obtained from a general relativistic simulation of the merger of two neutron stars with component masses of $1.3\,M_{\odot}$ and $1.4\,M_{\odot}$ [72], specifically, from the 7 ms post-merger snapshot shown on Fig. 1 (see [48] for a discussion of FFIs in this snapshot). These points (“Point A” and “Point B”) are chosen to illustrate two cases where a different branch of the CFI is unstable, namely, the minus (plus) mode for Point A (B). We detail the relevant parameters in Table 1. The opacities appearing in Eq. (1) are the absorption opacities obtained from NuLib [73] assuming the SFHo equation of state (like in the NSM simulation the data is extracted from). As discussed before, we assume that the values $N_{aa}$, $\overline{N}_{aa}$ in Table 1 correspond to the classical steady-state [^2] This classical state is not a genuine chemical equilibrium (for which $N_{xx}$ would be between $N_{ee}$ and $\overline{N}_{ee}$), but we effectively describe the escape of $\nu_{x}$ through the value $N_{xx}^{(\mathrm{cl})}<N_{ee}^{(\mathrm{cl})},\overline{N}_{ee}^{(\mathrm{cl})}$. $N_{aa}^{(\mathrm{cl})}$, $\overline{N}_{aa}^{(\mathrm{cl})}$, which is our initial condition.
Figure 1: Poloidal slice from the 7 ms post-merger snapshot of the “M1-NuLib” NSM simulation in [72]. The dashed line is the contour of $\nu_{e}$ flux factor equal to $0.1$.
@@ III.2 Numerical results
For Point A, the first $100\,\mathrm{\mu s}$ correspond to the growth of the unstable minus mode (exponential growth of $|N_{ex}|$ and $|\overline{N}_{ex}|$, see middle panel), which leads to flavor conversion $\nu_{e},\bar{\nu}_{e}\to\nu_{x}$. After the instability saturates, the densities of $\nu_{e}$ and $\bar{\nu}_{e}$ are repopulated very close to their classical equilibrium values, and oscillations have resulted in a net increase of $N_{xx}$. At the end of this first phase, $N_{xx}\simeq N_{ee}$. As can be seen in the bottom panel, this configuration is stable, consistent with Eq. (7a) with $\mathfrak{g}\simeq 0$. However, on a timescale $\propto\kappa_{x}^{-1}$, $N_{xx}$ is drawn back toward $N_{xx}^{(\mathrm{cl})}$, which brings the system into the unstable domain: the off-diagonal components grow again, and subsequent flavor conversion occurs, albeit with a smaller amplitude. After several cycles, the system converges toward a “quantum” equilibrium, where in particular $N_{xx}\neq N_{xx}^{(\mathrm{cl})}$ and flavor coherence is nonzero (with $|N_{ex}|/\mathrm{Tr}(N)$ at the percent level). For Point B, the conclusions are identical, reversing the roles of $\nu_{e}$ and $\bar{\nu}_{e}$, and the minus/plus modes. The difference of instability branches between A and B is visible in the evolutions of $N_{ex}$ and $\overline{N}_{ex}$, whose different behavior is consistent with the difference between $Q_{-}$ and $Q_{+}$ [see Eq. (7)].

### v3

@@ Abstract
Collisional flavor instabilities, driven by differing neutrino and antineutrino reaction rates, are expected to occur in dense astrophysical environments like supernovae and neutron star mergers, but have yet to be incorporated in large-scale simulations. We derive analytical expressions for the asymptotic state resulting from a homogeneous and isotropic instability, and apply these predictions to two representative conditions from a neutron star merger simulation. We emphasize the importance of using a collision term that allows for both damping of flavor coherence and relaxation back to the classical steady state. When this classical configuration is collisional-unstable, the resulting asymptotic state reflects a compromise between classical relaxation and flavor conversion, defining a “quantum” equilibrium with nonzero coherence. This analysis highlights the possibility of a tradeoff between classical and quantum effects, an important feature with regard to the inclusion of flavor oscillation physics into global simulations.
@@ I Introduction
In our current era of multimessenger astronomy, significant efforts are underway to model dense astrophysical objects such as core-collapse supernovae (CCSNe) or binary neutron star mergers (NSMs), with large-scale simulations incorporating a wide range of physics, including (magneto)hydrodynamics, general relativity, or the nuclear equation of state (see [1, 2, 3, 4, 5, 6] and references therein). A key ingredient in all simulations is neutrino transport [7, 8, 9, 10] yet to date, neutrino oscillations are not systematically taken into account. Over the last decades, however, various flavor transformation mechanisms have been identified, which are expected to occur in these environments and may have important consequences on the dynamics (see, e.g., [11, 12, 13, 14, 15, 16, 17, 18] for reviews).
@@
A direct approach, solving the “quantum” version of Boltzmann’s equation describing neutrino transport with flavor mixing—the quantum kinetic equation (QKE)—is made unfeasible by the vast differences of length scales and timescales involved. Nevertheless, to assess the potential consequences of neutrino flavor oscillations, simplified models have been used in CCSNe [19, 20, 21, 22], postmerger accretion disks [23, 24, 25], and recently, a full NSM simulation [26]. Other techniques, designed to mitigate the computational cost of a full multiangle QKE treatment, include an attenuation of the QKE Hamiltonian [27, 28, 29, 30, 31, 32], or the use of angular moments of the neutrino density matrix [33, 34, 35, 36, 37, 38, 39]. A convenient approach that can be embedded in classical simulations consists in designing a subgrid model of flavor transformation [40], which requires one to determine the state reached by the system post-flavor conversion. This has been heavily studied for fast flavor instabilities (FFIs), establishing that the final quasisteady state is characterized by some cancellation of the angular crossing between (anti)neutrino distributions, which can be predicted in various ways [41, 42, 43, 44, 45, 46, 47, 48, 49]. Such prescriptions have recently been used in global simulations of a postmerger disk [50] and CCSNe [51].
@@
A similar analytical prediction of the flavor conversion outcome remains, so far, unrealized for “collisional” flavor instabilities (CFIs), first evidenced in [52]. CFIs take place as neutrino self-interactions combine with different neutrino/antineutrino reaction rates to amplify flavor coherence, contrary to the standard expectation that collisions damp flavor coherence. Linear stability analysis [53, 54, 55, 56, 57] for a homogeneous and isotropic background has shown the existence of different collisional-unstable modes, with resonancelike and non-resonancelike regimes. The occurrence of CFIs has been investigated in actual environments, namely CCSNe [28, 58, 59, 60, 61, 62] and NSMs [53, 36, 63]. However, most studies consider a simplified collision term which only acts on flavor coherence. This still leads to a rich phenomenology, notably with multienergy neutrino gases, from flavor equipartition to flavor swaps [54, 55, 64, 65, 66]. Including the flavor-diagonal parts of the collision term is nevertheless important, as they lead for instance to a flavor swap in the resonancelike case [67], instead of flavor equipartition.
@@
This paper is structured as follows. In Sec. II, we present our model of neutrino evolution, which includes both classical relaxation and coherence damping components of the collision term. We also present the associated linear stability results, with details in Appendix A. In Sec. III, we numerically solve the evolution equations for configurations taken from a classical NSM simulation (see Appendix B for a convergence study), and derive analytical predictions of the asymptotic state that agree with the nonlinear outcome of collisional flavor conversion. These results show significant differences with earlier studies that use different implementations of the collision term, which is examined in Sec. IV. We summarize our findings and discuss the limitations and prospects of this study in Sec. V.
@@ II.1 Model and assumptions
where $\omega\equiv\Delta m^{2}/2E_{\nu}$ with $E_{\nu}$ the common neutrino energy, and we use the compact notation $c_{2\theta}=\cos(2\theta)$ and $s_{2\theta}=\sin(2\theta)$, with $\theta$ the mixing angle. In this work, we take $\Delta m^{2}=7.5\times 10^{-5}\,\mathrm{eV^{2}}$ and $\sin^{2}(\theta)=0.307$ (values corresponding to $\Delta m^{2}_{21}$ and $\theta_{12}$ in the 3-flavor case [71]). The matter term reads, also in the flavor basis,
@@
The last terms in Eq. (1) are an approximate description of collisions, using a “relaxation-time” approximation. The quantities $N^{(\mathrm{cl})}$, $\overline{N}^{(\mathrm{cl})}$ are flavor-diagonal and describe the classical steady-state of the system. The timescales associated to this relaxation are given by the opacities $\kappa_{e}$, $\bar{\kappa}_{e}$ and $\kappa_{x}=\bar{\kappa}_{x}$ (we assume, consistently with many large-scale simulations, identical evolutions for $\nu_{x}$ and $\bar{\nu}_{x}$). For $\nu_{e}$ and $\bar{\nu}_{e}$, this corresponds to emission/absorption processes, with absorption opacity $\kappa_{e}$ and associated emissivity $\kappa_{e}N_{ee}^{(\mathrm{cl})}$ (likewise for $\bar{\nu}_{e}$), but the treatment of heavy-lepton flavor (anti)neutrinos is more subtle. Indeed, it was shown in [61] that a system in beta- and pair-equilibrium is stable to CFIs. In an actual system that is unstable to CFI, the heavy-lepton flavor neutrinos $\nu_{x}$ deviate from this thermal equilibrium as they are not perfectly trapped. The classical steady-state $N_{xx}^{(\mathrm{cl})}$ is thus the result of a competition between emission/absorption and advection. An effective homogeneous description of the system can be done in a relaxation-time approximation, assuming the same form of a collision-like term as for electron flavor (anti)neutrinos but with an effectively reduced emissivity (see Sec. III.1 for an actual example), which leads to the form (1) of the evolution equation. In the following, we conflate “collisions” with “classical steady-state relaxation,” bearing in mind that this remains a limitation of our homogeneous description. We note that this description would also apply to a perfectly homogeneous and isotropic system with a nonstandard classical equilibrium. Finally, since we consider isotropic neutrino distributions, the scattering processes can be discarded.
We will assume that initially, the system is in the classical steady-state $N(t=0)=N^{(\mathrm{cl})}\equiv\mathrm{diag}\left(N_{ee}^{(\mathrm{cl})},N_{xx}^{(\mathrm{cl})}\right)$, and likewise for antineutrinos. This assumption allows us to focus on the competing effects of flavor conversion and classical relaxation. If the initial configuration was not a classical steady-state, the diagonal components would evolve on the collisional timescale while flavor coherence grows, but accounting for these effects would notably require to describe the advection of $\nu_{x}$. We thus focus on an idealized situation that allows for an analytical treatment, and provides guidance for future studies on larger spatial scales.
@@ II.2 Linear regime
We first establish the main features of the CFI in the linear regime, where $\lvert N_{ex}\rvert\ll 1$. Following [54], we write the $ex$ components of the QKEs as
$\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}N_{ex}\\ \overline{N}_{ex}\end{bmatrix}\simeq\frac{\omega s_{2\theta}}{2\sqrt{2}G_{F}}\begin{bmatrix}-\mathfrak{g}\\ \bar{\mathfrak{g}}\end{bmatrix}+\Lambda\begin{bmatrix}N_{ex}\\ \overline{N}_{ex}\end{bmatrix}\,,$ (4)
@@
$\Lambda=\begin{bmatrix}-\omega c_{2\theta}+\lambda-\bar{\mathfrak{g}}-\mathrm{i}\Gamma&\mathfrak{g}\\ -\bar{\mathfrak{g}}&\omega c_{2\theta}+\lambda+\mathfrak{g}-\mathrm{i}\overline{\Gamma}\end{bmatrix}\,,$ (5)
@@
$\mathrm{Im}\left(\Omega_{-}\right)$ $=\frac{\overline{\Gamma}\bar{\mathfrak{g}}-\Gamma\mathfrak{g}}{\mathfrak{g}-\bar{\mathfrak{g}}}\,,$ $Q_{-}$ $\propto\begin{bmatrix}\mathfrak{g}\\ \bar{\mathfrak{g}}\end{bmatrix}\,,$ (7a)
$\mathrm{Im}\left(\Omega_{+}\right)$ $=\frac{\Gamma\bar{\mathfrak{g}}-\overline{\Gamma}\mathfrak{g}}{\mathfrak{g}-\bar{\mathfrak{g}}}\,,$ $Q_{+}$ $\propto\begin{bmatrix}1\\ 1\end{bmatrix}\,.$ (7b)
We assumed that we are outside the resonancelike regime, which is the case in the examples studied below. The above derivation is valid for the initial configuration where $N_{ex}\simeq 0$ and $N_{aa}=N_{aa}^{(\mathrm{cl})}$. Specifically, Eq. (7) with $\mathfrak{g}=\mathfrak{g}^{(\mathrm{cl})}$ indicates whether this initial configuration is (un)stable. More generally, the linear stability results (7) are applicable at all times where $|N_{ex}(t)|\ll 1$, where they give in particular the instantaneous instability growth rate.
@@ III.1 NSM examples
We consider two configurations obtained from a general relativistic simulation of the merger of two neutron stars with component masses of $1.3\,M_{\odot}$ and $1.4\,M_{\odot}$ [72], specifically, from the 7 ms postmerger snapshot shown on Fig. 1 (see [48] for a discussion of FFIs in this snapshot). These points (“Point A” and “Point B”) are chosen to illustrate two cases where a different branch of the CFI is unstable, namely, the minus (plus) mode for Point A (B). We detail the relevant parameters in Table 1. The opacities appearing in Eq. (1) are the absorption opacities obtained from NuLib [73] assuming the SFHo equation of state (like in the NSM simulation the data is extracted from). As discussed before, we assume that the values $N_{aa}$, $\overline{N}_{aa}$ in Table 1 correspond to the classical steady-state [^2] This classical state is not a genuine chemical equilibrium (for which $N_{xx}$ would be between $N_{ee}$ and $\overline{N}_{ee}$), but we effectively describe the escape of $\nu_{x}$ through the value $N_{xx}^{(\mathrm{cl})}<N_{ee}^{(\mathrm{cl})},\overline{N}_{ee}^{(\mathrm{cl})}$. $N_{aa}^{(\mathrm{cl})}$, $\overline{N}_{aa}^{(\mathrm{cl})}$, which is our initial condition.
Figure 1: Poloidal slice from the 7 ms postmerger snapshot of the “M1-NuLib” NSM simulation in [72]. The dashed line is the contour of $\nu_{e}$ flux factor equal to $0.1$.
@@ III.2 Numerical results
For Point A, the first $100\,\mathrm{\mu s}$ correspond to the growth of the unstable minus mode (exponential growth of $|N_{ex}|$ and $|\overline{N}_{ex}|$, see middle panel), which leads to flavor conversion $\nu_{e},\bar{\nu}_{e}\to\nu_{x}$. After the instability saturates, the densities of $\nu_{e}$ and $\bar{\nu}_{e}$ are repopulated very close to their classical equilibrium values, and oscillations have resulted in a net increase of $N_{xx}$. At the end of this first phase, $N_{xx}\simeq N_{ee}$. As can be seen in the bottom panel, this configuration is stable, consistent with Eq. (7a) with $\mathfrak{g}\simeq 0$. However, on a timescale $\propto\kappa_{x}^{-1}$, $N_{xx}$ is drawn back toward $N_{xx}^{(\mathrm{cl})}$, which brings the system into the unstable domain; the off-diagonal components grow again, and subsequent flavor conversion occurs, albeit with a smaller amplitude. After several cycles, the system converges toward a “quantum” equilibrium, where in particular $N_{xx}\neq N_{xx}^{(\mathrm{cl})}$ and flavor coherence is nonzero (with $|N_{ex}|/\mathrm{Tr}(N)$ at the percent level). For Point B, the conclusions are identical, reversing the roles of $\nu_{e}$ and $\bar{\nu}_{e}$, and the minus/plus modes. The difference of instability branches between A and B is visible in the evolutions of $N_{ex}$ and $\overline{N}_{ex}$, whose different behavior is consistent with the difference between $Q_{-}$ and $Q_{+}$ [see Eq. (7)].